David bought three times as many hamburgers as rice toppings. Each Hamburger costs $25 and each rice topping is $75 more than each hamburger. If he spent $1050 in all, how many rice toppings did he buy?
Answer by 428225(90)

Let x = number of rice toppings bought
∴ 3x= number of hamburgers bought

You know that each hamburger cost $25 and that each rice topping was worth $75 dollars more than each hamburger, which means rice toppings cost $100.

∴ 3x(25) + x(100) = 1050

75x + 100x = 1050

175x = 1050

∴ x = 6

Therefore, David bought 6 rice toppings and 18 hamburgers.

To check the answer substitute 6 as x into the equation from earlier

3(6)(25) + 6(100) = 1050

450 + 600 = 1050

1050 = 1050

∴ Answer is correct.